OpenStudy (anonymous):

If the hypotenuse of a 45º-45º-90º triangle is 54 inches long, what is the length of one of the legs? Leave your answer in radical form.

OpenStudy (alana242):

Let x = the length of the legs of the right triangle 2x^2 = 54^2 x^2 = 54^2/2 x = 54/2^1/2 x = 54*sqrt of 2/2 x = 27*sqrt of 2